OpenStudy (swissgirl):

\( {(x,y) \in Z x Z: x+y = 10} \) Is this reflexive and transitive?

OpenStudy (swissgirl):

its only reflexive when x =5 so I am assuming that it isnt reflexive

OpenStudy (swissgirl):

I mean it doesnt seem transitive too me but what do u guys think?

OpenStudy (experimentx):

2 + 8 = 10 8 + 2 = 10

OpenStudy (helder_edwin):

u r right! it isn't reflexive

OpenStudy (swissgirl):

okkkkk

OpenStudy (swissgirl):

no like it cant be transitive

OpenStudy (helder_edwin):

it isn't transitive either.look at the example of @experimentX

OpenStudy (swissgirl):

righhtttt yaaa

OpenStudy (experimentx):

well ... solved!!

OpenStudy (swissgirl):

Great. Thanks guys :DDDDDDDDDDDDDD

OpenStudy (swissgirl):

so like r u sure its not reflexive?

OpenStudy (helder_edwin):

yes

OpenStudy (swissgirl):

ok

OpenStudy (anonymous):

It is not reflexive.

OpenStudy (helder_edwin):

the definition is \[ \large (\forall a\in\mathbb{Z})(aRa) \] but, as u pointed out, only 5 satisfies this

OpenStudy (swissgirl):

Thannkkksssss